For each of the following, find the slope, if it exists, of the line determined by the given pair of points:
a. (-4,1) and (5,2)
b. (-3,81) and (-3,198)
c. (1.0001,12) and (1,10)
d. (2, (sqrt)5) and (-2, and (sqrt)5)

Answer by Alan3354(69443)

For each of the following, find the slope, if it exists, of the line determined by the given pair of points:
a. (-4,1) and (5,2)
---
Slope m = diffy/diffx
m = (1-2)/(-4-5) = -1/-9
m = 1/9
=============
Do the others the same way.
b. (-3,81) and (-3,198)
c. (1.0001,12) and (1,10)
d. (2, (sqrt)5) and (-2, and (sqrt)5)
It's sqrt(5), not (sqrt)5